# Automated discovery and optimization of autonomous quantum error correction codes for a general open quantum system

### v1

@@ Abstract
We develop a method to search for the optimal code space, induced decay rates and control Hamiltonian to implement autonomous quantum error correction (AQEC) for a general open quantum system. The system is defined by a free-evolution Lindbladian superoperator, which contains the free Hamiltonian and naturally occurring decoherence terms, as well as the control superoperators. The performance metric for optimization is the fidelity between the projector onto the code space and the same projector after Lindbladian evolution for a specified time. We use a gradient-based search to update the code words, induced decay matrix and control Hamiltonian matrix. We apply the method to optimize AQEC codes for a variety of few-level systems. The four-level system with uniform decay rates offers a simple example for testing and illustrating the operation of our approach. The algorithm reliably succeeds in finding the optimal code in this case, while success becomes probabilistic for more complicated cases. For a five-level system with photon loss decay, the algorithm finds good AQEC codes, but these codes are not as good as the well-known binomial code. We use the binomial code as a starting point to search for the optimal code for a perturbed five-level system. In this case, the algorithm finds a code that is better than both the original binomial code and any other code obtained numerically when starting from a random initial guess.
@@ I Introduction
In the most commonly studied paradigm of QEC, the quantum system is allowed to evolve freely for some time, which results in a probabilistic leakage of the quantum state to parts of the Hilbert space outside the code space. A syndrome measurement then determines whether such an error has occurred. Importantly, it is necessary that neither the error process nor the syndrome measurement distinguish between the code words. A correction operation is subsequently performed, conditioned on the outcome of the syndrome measurement. The result of the measurement-correction sequence is that the system returns to its initial, possibly unknown, quantum state.
@@
There have been numerous studies on AQEC in the literature. Of particular interest to us are recent studies that used numerical methods to discover new AQEC protocols Foesel ; Wang . In another approach, the proposed physical system is designed to naturally achieve AQEC dynamics Kapit . There have also been studies on the fundamental mathematical conditions required to achieve AQEC Lebreuilly ; Lihm . A few experiments have demonstrated AQEC Gertler or closely related state-stabilization effects Leghtas in the past few years .
Similarly to conventional QEC protocols, AQEC protocols are typically designed based on our existing understanding and intuition about the open-system dynamics of a given quantum system, such as a qubit array or a harmonic oscillator. A good example of this situation is bosonic codes that are designed based on our understanding of the linear dynamics in a harmonic oscillator.
@@
In this work, we develop a method to perform a gradient-based numerical search to identify and optimize AQEC protocols for a general quantum system. The algorithm performs three search procedure to simultaneously find (1) the optimal encoding of the computational states 0 and 1 into quantum states of the available quantum system, (2) the induced decay rates that autonomously correct errors and (3) control Hamiltonian matrix elements that sometimes need to be added to the free Hamiltonian to complete the AQEC protocol. The algorithm works by trying to maximize the fidelity of an unknown quantum state with the same state after evolution under the effect of all the processes at play, naturally occurring and controllably induced, for a set amount of time. It should be noted that there have been related previous studies on the optimization of encoding, measurements and measurement-outcome-conditioned unitary operators in the standard measure-then-correct model of QEC Yamamoto2005 ; Yamamoto2007 ; Kosut ; Nautrup ; Su ; Olle , as well as the related topic of approximate QEC Leung ; Beny ; Ng ; Reimpell ; Fletcher2007 ; Fletcher2008 . Furthermore, we note that the philosophy of our approach is closely related to that of Ref. Wang , which focused on qubit-cavity systems.
@@
We analyze the performance of our numerical discovery and optimization approach for qudits with dimensions 4, 5 and 6. The algorithm consistently succeeds in finding the optimal codes in the simplest cases. The success in finding optimal codes becomes less certain as we go to more complicated decay models and/or higher dimensions. Nevertheless, the algorithm succeeds with some probability, which can be acceptable, since we can run the algorithm multiple times until we find a good solution, and we can use a good-but-not-optimal code to achieve AQEC. Furthermore, we demonstrate that the algorithm can be a powerful tool to optimize codes in cases when intuition allows us to make a good guess for an approximate AQEC code. This situation occurs naturally if the system parameters are close to a simple pattern in which a good AQEC code is known.
II Formulation of the autonomous QEC code design problem as an optimization problem
@@
$i\hbar\frac{d\rho}{dt}=\left[\hat{H},\rho\right]+\sum_{j}\left\{\hat{a}_{j}\rho\hat{a}_{j}^{\dagger}-\frac{1}{2}\hat{a}_{j}^{\dagger}\hat{a}_{j}\rho-\frac{1}{2}\rho\hat{a}_{j}^{\dagger}\hat{a}_{j}\right\},$ (1)
@@
When no action is taken to protect the quantum information, a quantum state experiences decoherence as described by Eq. (1) with the naturally occurring decay channels. To implement an AQEC protocol and protect quantum states from decoherence, we add appropriately chosen terms to the Lindblad equation, both in the Hamiltonian and in the form of additional, controllably induced decay channels:
$i\hbar\frac{d\rho}{dt}$ $=$ $\left[\left(\hat{H}+\sum_{k}\hat{O}_{k}\right),\rho\right]+$ (2)
@@
where $\hat{O}_{k}$ are Hamiltonian control operators, and $\hat{b}_{l}$ are controlled decay jump operators.
@@
where the operator, or superoperator, $\tilde{L}$ encodes all the information about the Hamiltonian and decay terms. The flattened density matrix after time $t$ is given by
@@
We are now ready to formulate the AQEC design problem as an optimization problem. The goal is to preserve any quantum superposition of the two code words as long as possible. This goal can be expressed in terms of aiming to implement a unitary operator, specifically the unit operator, which leaves any initial state unchanged. Since we are interested in protecting states that are in the code space at the initial time, we need an operator that acts as the unit operator on the code space but returns zero for any input state that is outside the code space, similarly to what was done in Ref. Ashhab2022 for optimizing two-qubit gates in larger Hilbert spaces. The operator that we need in this case is the projector on the code space:
$\hat{P}=\left|\tilde{0}\right\rangle\left\langle\tilde{0}\right|+\left|\tilde{1}\right\rangle\left\langle\tilde{1}\right|,$ (5)
@@
where $\left|\tilde{0}\right\rangle$ and $\left|\tilde{1}\right\rangle$ are the code words. We emphasize here that the projector $\hat{P}$ is also an $n^{2}\times n^{2}$ matrix, or superoperator, that operates on the flattened density matrix. As such, it has only four nonzero matrix elements when expressed in any basis that contains the code words as basis states. When applied to a density matrix, it keeps the matrix elements that correspond to the states $\left|\tilde{0}\right\rangle$, $\left|\tilde{1}\right\rangle$ and their superpositions, while all other parts of the density matrix are eliminated.
@@
In order to define a performance metric for the purpose of optimizing the AQEC code, we imagine that we let the system evolve for a finite duration, and we choose a value for the evolution time $\tau$, which can in principle be any finite value, i.e. $0<\tau<\infty$. We then define the fidelity for our computations as
$F=\frac{1}{4}{\rm Tr}\left\{\exp\left[\tilde{L}\tau\right]\hat{P}\right\}.$ (6)
@@
The factor $1/4$ is used because, in a basis containing the states $\left\{\left|\tilde{0}\right\rangle,\left|\tilde{1}\right\rangle\right\}$, the projector $\hat{P}$ has four matrix elements equal to one, with zeros everywhere else. The fidelity will therefore be equal to one when $\tau=0$, i.e. when the state does not change at all. Unless the system parameters allow perfect state preservation, all choices for the AQEC code will have $F<1$ for any nonzero value of $\tau$. It is worth noting that, since there is always a steady state which the system asymptotically approaches even in the presence of decoherence, there will always be AQEC codes with finite values of the fidelity, even in the limit $\tau\to\infty$. In our calculations, we generally set $\gamma\tau=1$, where $\gamma$ is the rate of the natural decay (i.e. error) process. We also note here that there is no unique definition of the fidelity for the purpose of searching for AQEC codes. In principle, any function that is equal to a certain value when the quantum state is preserved but decreases when we move away from perfect state protection will be an acceptable function in the calculations. We find the fidelity in Eq. (6) to be a computationally convenient choice, as it is obtained using standard matrix operations.
@@
Two points should be noted in relation to the fidelity and its use as a performance metric. Both points are also related to the fact that, when evaluating the performance of QEC codes, it seems intuitive to compare the effective lifetimes with and without the application of the QEC protocol, rather than analyzing a fidelity function. Firstly, in a typical AQEC setting, the system quickly reaches a near-steady state that contains a mixture of the code words and (a small population of) the error states. This near-steady state then slowly decays to the infinite-time, true steady state. As a result, as far as the long-time dynamics is concerned, the fidelity does not start its decay from the value $F=1$, but from a slightly reduced value. It is therefore in general not possible to translate a fidelity value into a decay rate. We can extract a decay rate by considering the fidelity as a function of time. But this task requires the evaluation of the fidelity at multiple time values and is computationally less straightforward. Secondly, the decay channels can be different with and without the application of the AQEC code. For example, in this work we focus on the case where the free evolution involves only one decoherence channel, namely energy relaxation. Considering that the system under the influence of the AQEC protocol is a driven dissipative system, the effective dynamics will in general be qualitatively different from the free-evolution dynamics. The decoheherence channels can be different, as we will show with at least one example in Sec. III. Furthermore, there can be multiple decoherence channels, each with its own rate, in the effective dynamics. As a result, it is not completely straightforward to make a comparison between the decoherence rates with and without the AQEC protocol based on individual fidelity values. When the decoherence rates between the two situations are different by orders of magnitude, the comparison between decoherence rates calculated from simple formulae seems to be logical. When the difference between the decoherence rates is small, say at the level of a factor of 2 or less, the comparison between rates can be misleading, depending on other details of the physical system under consideration. We therefore use the fidelity as a general purpose performance metric that is reasonably predictive and easy-to-calculate, while keeping in mind that any standard metric must be supplemented by a knowledge of the nature of the effective dynamics. We will also calculate ratios between the bare and effective lifetimes in some cases. As an estimate for the ratio between the effective decay rate under AQEC and the decay rate under free evolution, we will use the formula
@@
where $\beta_{1}$ and $\beta_{2}$ are chosen to give fidelities that are slightly below the initial value, such that the numerator gives a good approximation for the slope of the fidelity at early times. For realistic AQEC parameters, $\kappa$ should be largely insensitive to the exact values of $\beta_{1}$ and $\beta_{2}$, in addition to being insensitive to the fact that the fidelity starts at a value below $F=1$ even at $t=0$.
We use gradient-based techniques to identify good AQEC codes. As mentioned above, the AQEC code has three variable components: (1) the code space made up of the two code words, (2) the induced decay rates, and (3) the controlled Hamiltonian matrix elements. We use a single matrix $\hat{b}$ and a single matrix $\hat{O}$ for the second and third components of the code. We update all three components by evaluating the gradient of the fidelity with respect to the relevant parameters and shifting the AQEC code parameters in the direction of the gradient. In principle, we could calculate a unified gradient that incorporates the parameters of all three components of the AQEC code, but we divide the step of updating the code parameters into three distinct update steps, alternating between the three and updating each one of them in each optimization iteration. One reason to separate the update steps is that the code space has different mathematical properties from other two components of the code, as we will explain shortly. Furthermore, keeping the three update steps separate maintains conceptual simplicity in the algorithm.

### v2

@@ Abstract
We develop a method to search for the optimal code space, induced decay rates and control Hamiltonian to implement autonomous quantum error correction (AQEC) for a general open quantum system. The system is defined by a free-evolution Lindbladian superoperator, which contains the free Hamiltonian and naturally occurring decoherence terms, as well as the control superoperators. The performance metric for optimization in our algorithm is the fidelity between the projector onto the code space and the same projector after Lindbladian evolution for a specified time. We use a gradient-based search to update the code words, induced decay matrix and control Hamiltonian matrix. We apply our algorithm to optimize AQEC codes for a variety of few-level systems. The four-level system with uniform decay rates offers a simple example for testing and illustrating the operation of our approach. The algorithm reliably succeeds in finding the optimal code in this case, while success becomes probabilistic for more complicated cases. For a five-level system with photon loss decay, the algorithm finds good AQEC codes, but these codes are not as good as the well-known binomial code. We use the binomial code as a starting point to search for the optimal code for a perturbed five-level system. In this case, the algorithm finds a code that is better than both the original binomial code and any other code obtained numerically when starting from a random initial guess. Our results demonstrate the promise of using computational techniques to discover and optimize AQEC codes in future real-world quantum computers.
@@ I Introduction
In the most commonly studied paradigm of QEC, the quantum system is allowed to evolve freely for some time, which results in a probabilistic leakage of the quantum state to parts of the Hilbert space outside the code space. A syndrome measurement then determines whether such an error has occurred. Importantly, it is necessary that neither the error process nor the syndrome measurement distinguish between the code words. In other words, the syndrome measurement reveals for example whether an error has occurred or not, but it does not reveal whether the encoded qubit is in state 0 or 1. A correction operation is subsequently performed, conditioned on the outcome of the syndrome measurement. The result of the measurement-correction sequence is that the system returns to its initial, possibly unknown, quantum state.
@@
Similarly to conventional QEC protocols, AQEC protocols are typically designed based on our existing understanding and intuition about the open-system dynamics of a given quantum system, such as a qubit array or a harmonic oscillator. A good example of this situation is bosonic codes that are designed based on our understanding of the linear dynamics of a harmonic oscillator.
@@
In this work, we develop a method to perform a gradient-based numerical search to identify and optimize AQEC protocols for a general open quantum system. The algorithm performs three search procedures to simultaneously find (1) the optimal encoding of the computational states 0 and 1 into quantum states of the available quantum system, (2) the induced decay rates that autonomously correct errors and (3) control Hamiltonian matrix elements that sometimes need to be added to the free-evolution Hamiltonian to complete the AQEC protocol. The algorithm works by trying to maximize the fidelity of an unknown quantum state with the same state after evolution under the effect of all the processes at play, naturally occurring and controllably induced, for a set amount of time.
@@
We analyze the performance of our numerical discovery and optimization approach for qudits with dimensions 4, 5 and 6. The algorithm consistently succeeds in finding the optimal codes in the simplest cases. The success in finding optimal codes becomes less certain as we go to more complicated decay models and/or higher dimensions. Nevertheless, the algorithm succeeds with some probability, which can be acceptable, since we can run the algorithm multiple times until we find a good solution, and we can use a good-but-not-optimal code to achieve AQEC. Furthermore, we demonstrate that the algorithm can be a powerful tool to optimize codes in cases when intuition allows us to make a good guess for an approximate AQEC code. This situation occurs naturally if the system parameters deviate slightly from a simple pattern in which a good AQEC code is known.
II Related Recent work
There has been a significant amount of research activity on AQEC in recent years. Some studies investigated the fundamental mathematical conditions required to achieve AQEC Lihm ; Lebreuilly ; Kwon2025 , identifying analogues of the Knill-Laflamme conditions Knill for AQEC.
A few studies proposed realizing AQEC dynamics in specific, carefully designed systems, especially in superconducting circuits. In Ref. Kapit , Kapit proposed implementing AQEC in a sustem of two coupled superconducting qubits. This proposal was realized experimentally by Li et al. Li2024 . In Ref. Kwon2022 , Kwon et al. proposed realizing AQEC in a weakly anharmonic superconducting oscillator.
Other experiments include realization of a binomial code by Gertler et al. Gertler , a Gottesman-Kitaev-Preskill (GKP) code by Lachance-Quirion et al. LachanceQuirion , a qudit GKP code by Brock et al. Brock , all of which used 3-dimensional cavities coupled to superconducting qubits. A few years earlier, Leghtas et al. Leghtas had demonstrated state stabilization of two quantum states, although coherence between the two states was not preserved in that experiment. DeBry et al. DeBry and Li et al. Li2025 demonstrated AQEC in four-level codes in trapped ions.
Of particular interest to us are the recent theoretical studies that used numerical methods to discover new QEC and AQEC protocols Foesel ; Wang ; Zeng . Fösel et al. Foesel used reinforcement learning to find QEC codes. Wang et al. Wang used adjoint optimization to search for AQEC codes. We note that the philosophy of our approach is closely related to that of Ref. Wang . There are, however, clear differences between the two works. Wang et al. Wang focused on a harmonic oscillator coupled to one dissipative qubit for decay and one qubit for control, while we consider general open quantum systems. Zeng et al. Zeng used reinforcement learning to search for AQEC codes in a harmonic oscillator.
In the context of our work on the numerical discovery of AQEC codes, we should also mention the related work on the optimization of encoding, measurements and measurement-outcome-conditioned unitary operators in the standard measure-then-correct paradigm of QEC Reimpell ; Yamamoto2005 ; Fletcher2007 ; Yamamoto2007 ; Kosut ; Fletcher2008 ; PoulsenNautrup ; Bausch ; Olle ; Su ; Casanova . References Reimpell ; Yamamoto2005 ; Fletcher2007 ; Yamamoto2007 ; Kosut ; Fletcher2008 ; Olle addressed the question of whether various numerical optimization methods can efficiently find optimal QEC codes, including the code space and error correction operations. Poulsen Nautrup et al. PoulsenNautrup performed a numerical search that optimizes the connectivity graph in the surface code. Bausch et al. Bausch developed a machine-learning approach to achieve an optimal decoding strategy for noisy syndrome data when operating the surface code. References PoulsenNautrup ; Olle ; Su used reinforcement learning as the optimization tool for QEC code discovery. Casanova et al. Casanova use Riemannian optimization and included a performance metric that favours simple QEC codes over more complicated ones.
It is also worth mentioning the theoretical studies that investigated approximate QEC Reimpell ; Leung ; Beny ; Ng . In this case, the Knill-Laflamme conditions are not satisfied, and hence perfect QEC cannot be achieved. One can nevertheless implement protocols to prolong the lifetime of quantum information.
III Formulation of the autonomous QEC code design problem as an optimization problem
III.1 Lindblad equation of motion
@@
$\frac{d\rho}{dt}=-\frac{i}{\hbar}\left[\hat{H},\rho\right]+\sum_{j}\left\{\hat{a}_{j}\rho\hat{a}_{j}^{\dagger}-\frac{1}{2}\hat{a}_{j}^{\dagger}\hat{a}_{j}\rho-\frac{1}{2}\rho\hat{a}_{j}^{\dagger}\hat{a}_{j}\right\},$ (1)
@@
When no action is taken to protect the quantum information, a quantum state experiences decoherence as described by Eq. (1) with the naturally occurring decay channels. To implement an AQEC protocol and protect quantum states from decoherence, we add appropriately chosen terms to the Lindblad equation, both in the Hamiltonian and in additional terms that describe the controllably induced decay channels:
$\frac{d\rho}{dt}$ $=$ $-\frac{i}{\hbar}\left[\left(\hat{H}+\sum_{q}\hat{O}_{q}\right),\rho\right]+$ (2)
@@
where $\hat{O}_{q}$ are Hamiltonian control operators, and $\hat{b}_{l}$ are controlled decay jump operators.
@@
where the operator, or superoperator, $\tilde{L}$ encodes all the information about the Hamiltonian and decay terms. This rearrangement of the matrix elements simplifies the task of solving the differential equation and optimizing the control parameters in it. For example, the flattened density matrix after time $t$ is given by
@@
III.2 Defining target operation for optimal control calculation
@@
We are now ready to formulate the AQEC design problem as an optimization problem. We start by noting that we focus on encoding a single bit of quantum information, i.e. a Hilbert space with two quantum states, even though encoding larger Hilbert spaces is also possible. The goal is to preserve any quantum superposition of the two code words as long as possible. This goal can be expressed in terms of aiming to implement a unitary operator, specifically the unit operator, which leaves any initial state unchanged. Since we are interested in protecting states that are in the code space at the initial time, we need an operator that acts as the unit operator on the code space but is insensitive to any input state that is outside the code space, similarly to what was done in Ref. Ashhab2022 for optimizing two-qubit gates in larger Hilbert spaces. The operator that we need in this case is the projector on the code space:
[…]
@@
where $\left|\tilde{0}\right\rangle$ and $\left|\tilde{1}\right\rangle$ are the code words. We emphasize here that the projector $\hat{P}$ is also an $n^{2}\times n^{2}$ matrix, or superoperator, that operates on the flattened density matrix. As such, it has only four nonzero matrix elements when expressed in any basis that contains the code words as basis states. When applied to a density matrix, it keeps the matrix elements that correspond to the states $\left\{\left|\tilde{0}\right\rangle,\left|\tilde{1}\right\rangle\right\}$ and their superpositions, while all other parts of the density matrix are eliminated, i.e. they correspond to the eigenvalue zero of $\hat{P}$.
@@
In order to define a performance metric for the numerical optimization algorithm, we imagine that we let the system evolve for a finite duration, and we choose a value for the evolution time $\tau$, which can in principle be any finite value, i.e. $0<\tau<\infty$. We then define the fidelity between the implemented operation and the target operator $\hat{P}$ as
[…]
@@
This formula can be understood as follows: In a basis that contains the states $\left\{\left|\tilde{0}\right\rangle,\left|\tilde{1}\right\rangle\right\}$, each term inside the trace focuses on one of the four matrix elements of the density matrix in the code space and calculates how this term is degraded by the open-system dynamics described by $\tilde{L}$. The factor $1/4$ is used because the projector $\hat{P}$ has four matrix elements equal to one, with zeros everywhere else. The fidelity is equal to 1 if and only if the density matrix remains unaffected by the dynamical evolution, i.e. it experiences decoherence-free evolution. It should be noted that Eq. (6) treats the four matrix elements as if they were independent of each other, even though they are not. We find this formulation convenient when treating the density matrix as a one-dimensional vector, without worrying about the fact that this vector must obey the standard conditions for a physical density matrix. It should also be noted that there is no unique definition of the fidelity for the purpose of searching for AQEC codes. In principle, any function that is equal to a certain value when the quantum state is preserved but decreases when we move away from perfect state protection will be an acceptable function in the calculations. We find the fidelity in Eq. (6) to be a computationally convenient choice, as it is obtained using standard matrix operations.
III.3 Temporal evolution of the fidelity
The fidelity $F$ is obviously equal to one when $\tau=0$, i.e. when the quantum state has not started evolving. Unless the system parameters allow perfect state preservation, all choices for the AQEC code will have $F<1$ for any nonzero value of $\tau$. In our calculations, we generally set $\gamma\tau=1$, where $\gamma$ is the rate of the natural decay (i.e. error) process.
@@ III.3 Temporal evolution of the fidelity
Two points should be noted here in relation to the dynamics and the use of the fidelity as a performance metric. Both points are also related to the fact that, when evaluating the performance of QEC codes, it seems intuitive to compare the effective lifetimes with and without the application of the QEC protocol, rather than analyzing a fidelity function. Firstly, in a typical AQEC setting, the system quickly reaches a near-steady state that contains a mixture of the code words and (a small population of) the error states. This near-steady state then slowly decays to the infinite-time, true steady state. As a result, as far as the long-time dynamics is concerned, the fidelity does not start its decay from the value $F=1$, but from a slightly reduced value. It is therefore in general not straightforward to translate the fidelity value into a decay rate. We can extract a decay rate by considering the fidelity as a function of time. However, this task requires the evaluation of the fidelity at multiple time values, which would complicate the optimization algorithm. On the other hand, we can identify the optimal AQEC protocol by choosing a somewhat arbitrary value of $\tau$. Once we have obtained an AQEC code in this way, we can calculate the effective decay rate under the influence of the obtained protocol. Secondly, the decay channels can be different with and without the application of the AQEC protocol. For example, in this work we focus on the case where the free evolution involves only one decoherence channel, namely energy relaxation. Considering that the system under the influence of the AQEC protocol is a driven dissipative system, the effective dynamics will in general be qualitatively different from the free-evolution dynamics. The decoheherence channels can be different, as we will show with at least one example in Sec. IV. Furthermore, there can be multiple decoherence channels, each with its own rate, in the effective dynamics. As a result, it is not completely straightforward to make a comparison between the decoherence rates with and without the AQEC protocol based on individual fidelity values. When the decoherence rates between the two situations are different by orders of magnitude, the comparison between decoherence rates calculated from simple formulae seems to be logical. When the difference between the decoherence rates is small, say at the level of a factor of 2 or less, the comparison between rates can be misleading, depending on other details of the physical system under consideration. We therefore use the fidelity as a general purpose performance metric that is reasonably predictive and easy-to-calculate, while keeping in mind that any standard metric must be supplemented by a knowledge of the nature of the effective dynamics. We will also calculate ratios between the bare and effective lifetimes in some cases. As an estimate for the ratio between the effective decay rate under AQEC and the decay rate under free evolution, we will use the formula
@@
where $\beta_{1}$ and $\beta_{2}$ are chosen to give fidelities that are slightly below the initial value, such that the numerator gives a good approximation for the slope of the fidelity at early times. For realistic AQEC parameters, $\kappa$ should be largely insensitive to the exact values of $\beta_{1}$ and $\beta_{2}$, in addition to being insensitive to the fact that the fidelity starts at a value below $F=1$ even at $t$ values that are much smaller than the natural decay times.
III.4 Optimization algorithm
Figure 1: Flowchart that summarizes the algorithm. First, the three components of the AQEC code are initialized, possibly to random initial values. These three components are then updated by following the gradients to improve the fidelity defined in Eq. (6). After every iteration, the two termination condition are tested. When neither condition is satisfied, the AQEC code is updated further. When one of the termination conditions is satisfied, the algorithm outputs the optimized AQEC code.